# Merger History of Clustered Primordial Black Holes

### v2

@@ V Primordial black hole merger rate
where $\theta$ is Heaviside step function, $a_{\rm h}$ is the semimajor axis separating the hard and soft binaries defined by Eq. (16) and $j_{\rm cut}=1/2$ is the angular momentum above which the binaries are stable with respect to perturbations. The presented form of the suppression factor is physically explained in the previous section. Note that the influence of the initial clustering of PBHs on their merger rate was also considered in Refs. Bringmann et al. (2019); Young and Byrnes (2020); Atal et al. (2020); Crescimbeni et al. (2025), but without taking into account the perturbations of the binaries. We also note recent paper Holst et al. (2024) that studied the PBH mergers in clusters during the era of PBH domination in the very early Universe.
@@
[…]
@@
$T=\frac{3c^{5}\overline{x}^{4}\delta_{\rm PBH}^{4/3}}{170G^{3}m^{3}}\sim 10^{34}\left(\frac{m}{M_{\odot}}\right)^{-5/3}\,\rm{yr}$ (24)
@@
Figure 2: Time evolution of the merger rate for different values of the density contrast $\delta_{\rm PBH}$ and velocity dispersion in the cluster $\sigma=10$ km s-1. The purple dashed line shows the case of a cluster with $\sigma=10^{3}$ km s-1 and $\delta_{\rm PBH}=0.1$. The red dotted line shows the analytical estimate of the merger rate given by Eq. (V). The vertical dashed lines show the moment of time $t_{\rm sup}$, determined by Eq. (20), for $\delta_{\rm PBH}=2$ and $\delta_{\rm PBH}=1$, respectively.
To analyze the obtained result, we derive an approximate expression for the merger rate in the time region where the suppression factor is of the form $S=e^{-t/\tau(j,t)}$. Also for the beginning we will consider that the dispersion of velocities of black holes in the cluster is insignificant, i.e. the second term of gravitational focusing in the cross section (18) is the main one. As can be seen from Fig. (1), the main contribution to the integral (23) comes from the binaries with angular momenta $j\gtrsim j_{p}$. Numerical analysis shows that a good approximation is achieved by truncating the integral (23) from below at the value $j=0.82\,j_{p}$, and the suppression factor in the integration region can be neglected $e^{-t/\tau(j,t)}\rightarrow 1$. After the corresponding calculations we obtain
@@
×$\times\left(\frac{m}{M_{\odot}}\right)^{-27/14}\left(\frac{\sigma}{10\,\rm km\,s^{-1}}\right)^{29/14}.$ (26)
@@
This dependence is shown by the red dotted line in Fig. (2), from which we can see that the derived expression indeed approximates well the merger rate at $t>t_{\rm sup}$. Let us now consider the limiting case of a cluster with huge velocity dispersion. In this case, the cross section (17) becomes the geometric $\Sigma\propto a^{2}$. Using Eqs. (13) and (18) we obtain that the characteristic time of the perturbations scales as $\tau\propto t^{-1/2}j^{7/2}$. As in Sec. IV, determining the limiting angular momentum $\tau(t,j_{p})=t$ leads to the following relation $j_{p}\propto t^{3/7}$. Truncating the integral (23) from below at angular momentum $\sim j_{p}$ (similar to the derivation Eq. (V)) gives the time dependence of the merger rate $\mathcal{R}\propto t^{-61/28}$. Note that in this case the power-law decay of the merger rate will change to exponential over time when the condition $j_{p}(t)=1$ is reached, as follows from Eq. (22). This is due to the fact that binaries with a sufficiently long lifetime are soft in the cluster. That is, even pairs of PBHs in circular orbits (and even more so in elliptical orbits) will be destroyed due to perturbations, which is also seen in Fig. 3 in the region of $\delta_{\rm PBH}>1$ and $\sigma>100$ km s-1
@@
Let us now turn to the discussion of larger time scales, where the merger rate experiences a sharp change in slope and comes to a dependence $\mathcal{R}\propto t^{-5/8}$. This result is more of an artifact of our model rather than a physical effect. This behavior arises from the specific form of the suppression factor (22), where we assume that the orbital parameters of binaries with sufficiently large angular momenta $j>j_{\rm cut}$ do not change under perturbations. Then it is this region of angular momenta that will give the dominant contribution to the integral (23), so it is evaluated by a dimensionless constant. The time dependence arises only from the pre-integral factor, which explains the result. Moreover, changing the value of $j_{\rm cut}$ will only affect the amplitude of the merger rate, not its logarithmic slope. As noted in Section IV, a correct description requires taking into account the change in the distribution of the orbital parameters of the binaries under perturbations. Assuming that the distribution near large angular momenta becomes the power-law form $dP/dj\propto j^{k}$, the merger rate will vary as $\mathcal{R}\propto t^{(k-6)/7}$ Franciolini et al. (2022), with $k=1$ corresponding to the thermal distribution. Thus, in the physically realistic case, we expect the merger rate to also vary only slightly with time. However, the main result of this work is that there is a time window in which there is a sharp drop in the merger rate, which significantly diverges from the predictions of models in which the PBHs are initially distributed according to a Poisson law in space. This provides an opportunity to test the clustering of PBHs in future GW observations.
@@
Figure 3: Contour plot of the modern merger rate of PBH binaries $\mathcal{R}_{0}$ (in units of [Gpc-3 yr-1]) calculated for the PBH fraction $f=0.01$ for different clustering parameters $\delta_{\rm PBH}$ and the velocity dispersion $\sigma$ in the cluster, the PBH mass $m=10\,M_{\odot}$. Regions in parameter space incompatible with the LIGO-Virgo-KAGRA merger rate $\mathcal{R}_{0}=18-44$ Gpc-3 yr-1 Abbott et al. (2023a) are ruled out by GW observations. For other PBH fractions $f\neq 0.01$, the merger rate should be rescaled by a factor of $f/0.01$ (also the condition $f\leq\delta_{\rm PBH}$ should be satisfied).
@@
As noted above, perturbations of binaries with small eccentricities (almost circular orbits) make them harder, hence the merger time will decrease and therefore the merger rate of such binaries will increase. Therefore, our results in Fig.2 for $\delta_{\rm PBH}>1$ and time scales $\sim t_{0}$ (in this case, binaries with $j>j_{\rm cut}$ merge) should be considered as a conservative lower estimate. Figure  3 shows the contour plot of the modern merger rate as a function of the clustering parameter $\delta_{\rm PBH}$ and the velocity dispersion $\sigma$ in the cluster. For the case $f\simeq 0.01$ the majority of the clustering parameter space is consistent with current GW observation. Also one can see that there exists a region of parameters near $\delta_{\rm PBH}\gtrsim 1$, where one can explain all dark matter in terms of PBHs with mass $\sim 10\,M_{\odot}$. However, many other constraints imposed on the PBHs of these masses exclude this possibility. With a significant increase in the clustering parameter $\delta_{\rm PBH}\gg 1$, the merger rate is determined by binaries with nearly circular orbits. Moreover, increasing $\delta_{\rm PBH}$ leads to an enhancement of the modern merger rate, as can also be seen in Fig. 2. Thus, the early clustering effect makes it virtually impossible to relax the gravitational-wave constraints to the level where all dark matter could consist of PBHs. This conclusion is also consistent with the results obtained in Ref. Bringmann et al. (2019). However it should be noted, that in the region $\delta_{\rm PBH}\gtrsim 1$ and $\sigma\gtrsim 100$ km/s, the merger rate is exponentially suppressed (according to Eq. (22)), since in this case all binary systems are soft and undergo destruction. In the opposite limit $\delta_{\rm PBH}\ll 1$ clusters are formed at enough late stages, that significantly reduces the efficiency of perturbations on binary systems (see Eq. (20)), and we return to the situation of PBHs unclustered at birth, for which the standard constraints on their fraction in dark matter $f$ are valid.
@@ VI Stochastic gravitational wave background
In the previous section the behavior of the PBH merger rate on cosmological time scales was considered. However individual events of black hole mergers at high redshifts may remain observationally unresolved, nevertheless their total contribution to the energy density of gravitational waves in the Universe can be significant. In this section, we discuss the impact of our results on the characteristics of the GW background, which is traditionally described by the relative spectral energy density Phinney (2001)
@@
[…]
where $\nu_{s}=(1+z)\nu$ is the gravitational wave frequency in the source system and $dE_{\rm gw}/d\ln\nu_{s}$ is the gravitational wave energy emitted during the merger of a binary black hole in a logarithmic bin of frequency. In this paper we use the inspiral-merger-ringdown energy spectrum Ajith et al. (2008, 2011). In fact, Eq. (VI) is the sum of the energy released from all merger events summed over all redshifts. It should be noted that the analysis of the stochastic GW background from PBH mergers has also been studied in Refs. Mandic et al. (2016); Wang et al. (2018); Clesse and Garc´ıa-Bellido (2017b); Atal et al. (2020); Bavera et al. (2022); Mukherjee and Silk (2021); Garc´ıa-Bellido et al. (2022); Braglia et al. (2021); Atal et al. (2022).
@@
[…]
@@
Figure 4: Spectral energy density of the stochastic gravitational-wave background from mergers of PBHs with $m=10\,M_{\odot}$ mass for different values of the clustering parameter $\delta_{\rm PBH}$ and $\sigma=10$ km s-1. For illustration purposes we adopt $f=0.1$ here, though microlensing constraints typically favor $f\sim 0.01$. The dotted line shows the analytic approximation is given by Eq. (VI), with the minimum plotting frequency determined by $\nu_{\rm sup}$ from Eq. (32). The dash-dotted line line shows the case of the Poisson (with no initial clustering) distribution of PBHs normalized such that the modern merger rate $\mathcal{R}_{0}=10$ Gpc-3 yr-1. The shaded areas show the projected sensitivity of some future GW detectors, as well as the upper limit from the O3 run results Abbott et al. (2021).
@@
where, as before, we have used the dust stage approximation $t(z)=2/3H(z)$, since we are interested in the high redshift limit. Taking the maximum frequency of the GW in Eq. (VI) as $\nu_{\rm max}=0.057\,c^{3}/Gm$ (corresponding to the spectrum break $dE_{\rm gw}/d\ln{\nu_{s}}$ at higher frequencies Ajith et al. (2008)), then good agreement with numerical calculations is obtained with the estimate $A\sim 1$:
@@
×$\times\left(\frac{m}{M_{\odot}}\right)^{-13/4}\left(\frac{\nu}{10\,\rm{Hz}}\right)^{-65/28}.$ (31)
@@
Thus, in the frequency range $\nu>\nu_{\rm sup}$, the spectral density of gravitational waves is characterized by a negative slope $\Omega_{\rm gw}\propto\nu^{-65/28}$. At lower frequencies, the classical dependence $\Omega_{\rm gw}\propto\nu^{2/3}$, typical for merging black hole binaries. However, in the frequency band of space-based interferometers like LISA, a excess of $\Omega_{\rm gw}$ is expected compared to the case of PBHs with Poisson spatial distribution (and perhaps even over astrophysical sources). Note that if the PBH fraction $f\ll 0.1$, such features in the GW energy spectrum will be unobservable against other cosmological GW contributions. Let us now discuss the frequencies 100$\sim$ 100∼ 100 Hz. Provided that PBHs of tens solar masses do not dominate the observed GW events, then in these frequencies the main contribution to the GW background comes from mergers of black holes of stellar origin Mandic et al. (2016), which also leads to the standard dependence $\Omega_{\rm gw}\propto\nu^{2/3}$. However, a comprehensive analysis of the GW background taking into account various PBHs clustering scenarios and the contribution of astrophysical sources is beyond the scope of this paper and will be addressed in a separate article. Nevertheless, as can be seen from Fig. 4, there is parameter region in which future GW detectors will be able to test the possibility of PBHs clustering.

### v3

@@ V Primordial black hole merger rate
where $\theta$ is Heaviside step function, $a_{\rm h}$ is the semimajor axis separating the hard and soft binaries defined by Eq. (16) and $j_{\rm cut}=1/2$ is the angular momentum above which the binaries are stable with respect to perturbations. The presented form of the suppression factor is physically explained in the previous section. The limitations of a sharp transition at angular momentum $j_{p}$ in the suppression model (22) are presented at the end of this section. A more physically realistic case, which must take into account the complex dynamics of binary-single PBH interaction at large binary angular momentum $j\sim 1$, is discussed in the appendix A. Note that the influence of the initial clustering of PBHs on their merger rate was also considered in Refs. Bringmann et al. (2019); Young and Byrnes (2020); Atal et al. (2020); Crescimbeni et al. (2025), but without taking into account the perturbations of the binaries. We also note recent paper Holst et al. (2024) that studied the PBH mergers in clusters during the era of PBH domination in the very early Universe.
@@
$\frac{dP}{dt}=\frac{3}{16}\frac{\delta_{\rm PBH}^{2}}{T}\left(\frac{t}{T}\right)^{-5/8}\int\limits_{j_{\rm min}}^{1}j^{-37/8}S(t,j)\,dj,$ (23)
@@
$T=\frac{3c^{5}\overline{x}^{4}\delta_{\rm PBH}^{4/3}}{170G^{3}m^{3}}\sim 10^{34}\left(\frac{m}{M_{\odot}}\right)^{-5/3}\,\rm{yr},$ (24)
@@
Figure 2: Time evolution of the merger rate for different values of the density contrast $\delta_{\rm PBH}$ and velocity dispersion in the cluster $\sigma=10$ km s-1 is shown by the solid lines. The purple dot-dashed line shows the case of a cluster with $\sigma=10^{3}$ km s-1 and $\delta_{\rm PBH}=0.1$. The red dotted line shows the analytical estimate of the merger rate given by Eq. (V). The dashed purple line shows the PBH merger rate when the effects of binary perturbations are negligible (corresponding to the unclustered case). The vertical dashed lines show the moment of time $t_{\rm sup}$, determined by Eq. (20), for $\delta_{\rm PBH}=2$ and $\delta_{\rm PBH}=1$, respectively.
To analyze the obtained result, we derive an approximate expression for the merger rate in the time region where the suppression factor is of the form $S=e^{-t/\tau(j,t)}$. Also for the beginning we will consider that the velocity dispersion of PBHs in the cluster is insignificant, i.e. the second term of gravitational focusing in the cross section (18) is the leading one. As can be seen from Fig. (1), the main contribution to the integral (23) comes from the binaries with angular momenta $j\gtrsim j_{p}$. Numerical analysis shows that a good approximation is achieved by truncating the integral (23) from below at the value $j=0.82\,j_{p}$, and the suppression factor in the integration region can be neglected $e^{-t/\tau(j,t)}\rightarrow 1$. After the corresponding calculations we obtain
@@
$\times\left(\frac{m}{M_{\odot}}\right)^{-27/14}\left(\frac{\sigma}{10\,\rm km\,s^{-1}}\right)^{29/14}.$ (26)
@@
This dependence is shown by the red dotted line in Fig. (2), from which we can see that the derived expression indeed approximates well the merger rate at $t>t_{\rm sup}$. Let us now consider the limiting case of a cluster with huge velocity dispersion. In this case, the cross section (17) becomes the geometric $\Sigma\propto a^{2}$. Using Eqs. (13) and (18) we obtain that the characteristic time of the perturbations scales as $\tau\propto t^{-1/2}j^{7/2}$. As in Sec. IV, determining the limiting angular momentum $\tau(t,j_{p})=t$ leads to the following relation $j_{p}\propto t^{3/7}$. Truncating the integral (23) from below at angular momentum $\sim j_{p}$ (similar to the derivation Eq. (V)) gives the time dependence of the merger rate $\mathcal{R}\propto t^{-61/28}$. Note that in this case the power-law decay of the merger rate will change to exponential over time when the condition $j_{p}(t)=1$ is reached, as follows from Eq. (22). This is due to the fact that binaries with a sufficiently long lifetime are soft in the cluster. That is, even PBH binaries in circular orbits (and even more so in elliptical orbits) will be destroyed due to perturbations, which is also seen in Fig. 3 in the region of $\delta_{\rm PBH}>1$ and $\sigma>100$ km s-1
@@
Let us now turn to the discussion of larger time scales, where the merger rate experiences a sharp change in slope and comes to a dependence $\mathcal{R}\propto t^{-5/8}$. This result is more of an artifact of our model rather than a physical effect. This behavior arises from the specific form of the suppression factor (22), where we assume that the orbital parameters of binaries with sufficiently large angular momenta $j>j_{\rm cut}$ do not change under perturbations. Then it is this region of angular momenta that will give the dominant contribution to the integral (23), so it is evaluated by a dimensionless constant. The time dependence arises only from the pre-integral factor, which explains the result. Moreover, changing the value of $j_{\rm cut}$ will only affect the amplitude of the merger rate, not its logarithmic slope. As noted in Section IV, a correct description requires taking into account the change in the distribution of the orbital parameters of the binaries under perturbations. In Appendix A, we develop a simple physical model for the merger of these tight PBH binaries, taking into account the evolution of the distribution of their parameters $a$ and $j$ due to perturbations in clusters. In particular, assuming that the angular momentum distribution becomes the power-law form $dP/dj\propto j^{n}$, the merger rate will vary as $\mathcal{R}\propto t^{(n-6)/7}$ (see also Ref. Franciolini et al. (2022)), with $n=1$ corresponding to the thermal distribution. Thus, in the physically realistic case, we expect the merger rate to also vary only slightly with cosmological time. However, the main result of this work is that there is a time window in which there is a sharp drop in the merger rate, which significantly diverges from the predictions of models in which the PBHs are initially distributed according to a Poisson law in space. This provides an opportunity to test the clustering of PBHs in future GW observations.
@@
Figure 3: Contour plot of the current merger rate of PBH binaries $\mathcal{R}_{0}$ (in units of [Gpc-3 yr-1]) calculated for the PBH fraction $f=0.01$ for different clustering parameters $\delta_{\rm PBH}$ and the velocity dispersion $\sigma$ in the cluster, the PBH mass $m=10\,M_{\odot}$. Regions in parameter space incompatible with the LIGO-Virgo-KAGRA merger rate $\mathcal{R}_{0}=18-44$ Gpc-3 yr-1 Abbott et al. (2023a) are ruled out by GW observations. For other PBH fractions $f\neq 0.01$, the merger rate should be rescaled by a factor of $f/0.01$ (also the condition $f\leq\delta_{\rm PBH}$ should be satisfied).
@@
As noted above, perturbations of binaries with small eccentricities (almost circular orbits) make them harder, hence the merger time will decrease and therefore the merger rate of such binaries will increase. Therefore, our results in Fig.2 for $\delta_{\rm PBH}>1$ and time scales $\sim t_{0}$ (in this case, binaries with $j>j_{\rm cut}$ merge) should be considered as a conservative lower estimate. Figure 3 shows the contour plot of the current merger rate as a function of the clustering parameter $\delta_{\rm PBH}$ and the velocity dispersion $\sigma$ in the cluster. For the case $f\simeq 0.01$ the majority of the clustering parameter space is consistent with current GW observation. Also one can see that there exists a region of parameters near $\delta_{\rm PBH}\gtrsim 1$, where one can explain all dark matter in terms of PBHs with mass $\sim 10\,M_{\odot}$. However, many other constraints imposed on the PBHs of these masses exclude this possibility. With a significant increase in the clustering parameter $\delta_{\rm PBH}\gg 1$, the merger rate is determined by binaries with nearly circular orbits. Moreover, increasing $\delta_{\rm PBH}$ leads to an enhancement of the current merger rate, as can also be seen in Fig. 2. Thus, the early clustering effect makes it virtually impossible to relax the GW constraints to the level where all dark matter could consist of PBHs. This conclusion is also consistent with the results obtained in Ref. Bringmann et al. (2019). However it should be noted, that in the region $\delta_{\rm PBH}\gtrsim 1$ and $\sigma\gtrsim 100$ km/s, the merger rate is exponentially suppressed (according to Eq. (22)), since in this case all binary systems are soft and undergo destruction. In the opposite limit $\delta_{\rm PBH}\ll 1$ clusters are formed at enough late stages, that significantly reduces the efficiency of perturbations on binary systems (see Eq. (20)), and we return to the situation of PBHs unclustered at birth, for which the standard constraints on their fraction in dark matter $f$ are valid.
@@ VI Stochastic gravitational wave background
In the previous section the behavior of the PBH merger rate on cosmological time scales was considered. However individual events of black hole mergers at high redshifts may remain observationally unresolved, nevertheless their total contribution to the energy density of GW in the Universe can be significant. In this section, we discuss the impact of our results on the characteristics of the GW background, which is traditionally described by the relative spectral energy density Phinney (2001)
@@
$=\frac{1}{\rho_{\rm crit}c^{2}}\int_{0}^{\infty}dz\,\frac{\mathcal{R}(z)}{H(z)(1+z)^{2}}\frac{dE_{\rm gw}}{d\ln\nu_{s}},$ (27)
where $\nu_{s}=(1+z)\nu$ is the GW frequency in the source system and $dE_{\rm gw}/d\ln\nu_{s}$ is the GW energy emitted during the merger of a binary black hole in a logarithmic bin of frequency. In this paper we use the inspiral-merger-ringdown energy spectrum Ajith et al. (2008, 2011). In fact, Eq. (VI) is the sum of the energy released from all merger events summed over all redshifts. It should be noted that the analysis of the stochastic GW background from PBH mergers has also been studied in Refs. Mandic et al. (2016); Wang et al. (2018); Clesse and García-Bellido (2017b); Atal et al. (2020); Bavera et al. (2022); Mukherjee and Silk (2021); García-Bellido et al. (2022); Braglia et al. (2021); Atal et al. (2022).
@@
$\Omega_{\rm gw}=A\nu^{2/3}\frac{\mathcal{R}_{0}\pi^{2/3}(Gm)^{5/3}}{3\sqrt{2}G\rho_{\rm crit}c^{2}}\int\limits_{0}^{(\nu_{\rm max}/\nu)-1}\frac{(1+z)^{\kappa+2/3}}{(1+z)^{2}H(z)}\,dz$
@@
Figure 4: Spectral energy density of the stochastic GW background from mergers of PBHs with $m=10\,M_{\odot}$ mass for different values of the clustering parameter $\delta_{\rm PBH}$ and $\sigma=10$ km s-1. For illustration purposes we adopt $f=0.1$ here, though microlensing constraints typically favor $f\sim 0.01$. The dashed line shows the analytic approximation is given by Eq. (VI), with the minimum plotting frequency determined by $\nu_{\rm sup}$ from Eq. (32). The dash-dotted line line shows the case of the Poisson (with no initial clustering) distribution of PBHs normalized such that the current merger rate $\mathcal{R}_{\rm PBH}(t_{0})=10$ Gpc-3 yr-1 and changes with redshift as $\mathcal{R}\propto t^{-34/37}$. The red dotted line shows the case of astrophysical black holes (see Appendix B for details) and $\mathcal{R}_{\rm ABH}(t_{0})=40$ Gpc-3 yr-1. The shaded areas show the projected sensitivity of some future GW detectors, as well as the upper limit from the O3 run results Abbott et al. (2021).
@@
where, as before, we have used the dust stage approximation $t(z)=2/3H(z)$, since we are interested in the high redshift limit. Taking the maximum frequency of the GW in Eq. (VI) as $\nu_{\rm max}=0.057\,c^{3}/Gm$ (corresponding to the spectrum break $dE_{\rm gw}/d\ln{\nu_{s}}$ at higher frequencies Ajith et al. (2008)), that approximation means that the inspiral phase is formally extended to the ringdown frequency. Good agreement with numerical calculations is then obtained with the estimate $A\sim 1$:
@@
$\times\left(\frac{m}{M_{\odot}}\right)^{-13/4}\left(\frac{\nu}{10\,\rm{Hz}}\right)^{-65/28}.$ (31)
@@
Thus, in the frequency range $\nu>\nu_{\rm sup}$, the spectral density of GW is characterized by a negative slope $\Omega_{\rm gw}\propto\nu^{-65/28}$. At lower frequencies, the classical dependence $\Omega_{\rm gw}\propto\nu^{2/3}$, typical for merging black hole binaries. However, in the frequency band of space-based interferometers like LISA, a excess of $\Omega_{\rm gw}$ is expected compared to the case of PBHs with Poisson spatial distribution and even astrophysical black hole mergers are shown by the red dotted line in Fig. 4. Thus, the detection of strong GW background in this frequency range, coupled with the absence of a corresponding signal on ground-based detectors (when interpolated to higher frequencies $\Omega_{\rm gw}\propto\nu^{2/3}$) will hint at the presence of PBHs, which must be highly clustered. However note that if the PBH fraction $f\ll 0.1$, such features in the GW energy spectrum will be unobservable against other cosmological GW contributions. Let us now discuss the frequencies $\sim 100$ Hz. Provided that PBHs of tens solar masses do not dominate the observed GW events, then at these frequencies the main contribution to the GW background comes from mergers of black holes of stellar origin Mandic et al. (2016), which also leads to the standard dependence $\Omega_{\rm gw}\propto\nu^{2/3}$, as can be seen in Fig. 4. However, a comprehensive analysis of the GW background taking into account various PBHs clustering scenarios and the contribution of astrophysical sources is beyond the scope of this paper and will be addressed in a separate article. Nevertheless, as can be seen from Fig. 4, there is parameter region in which future GW detectors will be able to test the possibility of PBHs clustering.
Table 1: Scaling of the merger rate $\mathcal{R}$ and stochastic GW background $\Omega_{\rm gw}$ for different cases of black hole mergers: with and without initial clustering for PBHs, and astrophysical black holes. For the clustered PBH scenario, the scaling for $\mathcal{R}$ and $\Omega_{\rm gw}$ applies after the suppression time $t_{\rm sup}$ and frequency $\nu_{\rm sup}$, respectively (see Eqs. (20) and (32)). Here, we introduce the notation $\sigma_{10}$ is the velocity dispersion in units of 1010 km s-1 and $m_{1}$ is the PBH mass in $M_{\odot}$, in order to write the expression for $t_{\rm sup}$.
PBH clustering with low $\sigma$ PBH clustering with high $\sigma$ PBH without clustering ABH
$\mathcal{R}$ $\propto t^{-45/14}$ $\propto t^{-61/28}$ $\propto t^{-34/37}$ see Fig. 6
$\Omega_{\rm gw}$ $\propto\nu^{-65/28}$ $\propto\nu^{-43/56}$ $\propto\nu^{2/3}$ $\propto\nu^{2/3}$
$t_{\rm sup}$ [Myr] $4.1\,\delta_{\rm PBH}^{-120/41}\sigma_{10}^{37/41}m_{1}^{-19/41}$ $5.9\,\delta_{\rm PBH}^{-92/45}\sigma_{10}^{-37/45}m_{1}^{-1/45}$
$\nu_{\rm sup}$ [Hz] $2.2\,\delta_{\rm PBH}^{-80/41}\sigma_{10}^{74/123}m_{1}^{-161/123}$ $2.8\,\delta_{\rm PBH}^{-184/135}\sigma_{10}^{-74/135}m_{1}^{-2/135}$